# Rich Semantic Knowledge Enhanced Large Language Models for Few-shot Chinese Spell Checking

### v1

@@ Abstract
Chinese Spell Checking (CSC) is a widely used technology, which plays a vital role in speech to text (STT) and optical character recognition (OCR). Most of the existing CSC approaches relying on BERT architecture achieve excellent performance. However, limited by the scale of the foundation model, BERT-based method does not work well in few-shot scenarios, showing certain limitations in practical applications. In this paper, we explore using an in-context learning method named RS-LLM (Rich Semantic based LLMs) to introduce large language models (LLMs) as the foundation model. Besides, we study the impact of introducing various Chinese rich semantic information in our framework. We found that by introducing a small number of specific Chinese rich semantic structures, LLMs achieve better performance than the BERT-based model on few-shot CSC task. Furthermore, we conduct experiments on multiple datasets, and the experimental results verified the superiority of our proposed framework.
@@
[…]
@@ 1 Introduction
As an ideogram, the usage and structure of Chinese are very different from English, which leads to different challenges for CSC and SC. First of all, the pronunciation of Chinese varies greatly, and it is difficult to easily infer the glyphs from the pronunciation. When listening to a piece of Chinese speech, if you do not understand the context, the text result obtained based on the speech is likely to contain a large number of homophones. In addition, the glyph structure of Chinese is more diverse, resulting in more types of errors. Therefore, CSC task mainly needs to handle two types of error texts. One is a text based on the STT system, which contains a large number of homophonic errors. The other is text generated by OCR-based systems that mainly contains glyph errors. To address these two types of spelling errors, most of the existing studies use models based on BERT architecture, then introducing the external glyph-phonetic features Ji et al. (2021); Xu et al. (2021); Ji et al. (2021); Li et al. (2022).
@@ 3.2 Task Definition
Formulate definition of CSC: CSC focuses on identifying and rectifying spelling errors in a given text sequence. Essentially, it involves processing a sequence of characters $X=\{x_{1},x_{2},...,x_{n}\}$ and producing a corrected sequence $Y=\{y_{1},y_{2},...,y_{n}\}$, where $n$ is the number of words. Here, $X$ denotes the initial text with potential errors, and $Y$ signifies the amended correct text. The two sequences $X$ and $Y$ have the same length.
@@
3.3 Challenges of LLMs based CSC Tasks
The versatility of LLMs gives them significant text polishing capabilities. Since there are no specific restrictions on input prompts, LLMs tend to perform freely in CSC tasks. However, free play of LLMs may result in LLMs outputting sentences that are completely grammatically correct. However, the sentences output by LLMs are different from the standards established by existing CSC datasets and evaluation indicators. Therefore, these existing traditional datasets are used without specific restrictions on input prompts. It becomes challenging to objectively and realistically evaluate the spell checking performance of LLMs. An example of using LLM (ChatGPT) to perform a CSC task without setting specific restrictions on input prompts is shown in Fig.1. From Fig.1 it can be observed that two main problems are prone to occur in LLM when performing CSC tasks, one is the length of the input sentence and the output sentence are inconsistent. For instance, the sentence "I am really bald (我真胡秃)", which should be corrected to "I am really confused (我真糊涂)". However, LLM correct this sentence to "I am really sorry (我真的抱歉)". Another problem is that LLM easily rewrites the input sentences semantically. For example, "I hope you have a good time dancing without (希望你们好好的跳无)", in which "dancing without (跳无)" should be corrected to "dancing (跳舞)". Instead, LLM correct this sentence to "I hope you have a good New Year (希望你们好好的过年)".
Figure 2: Task-specific few-shot prompt for CSC tasks. We marked the semantic structure information (speech and radicals) and key information related to the task features in the prompts in different color fonts.
@@ 4.1 Chinese Rich Semantic Corpus
Chinese, a logo-graphic language, inherently possesses a rich semantic depth in its character glyph, potentially enhancing the expressiveness of LLMs. Our research focuses on the GB2312 simplified Chinese coding table, a standard set by China’s State Administration of Standards on May 1, 1981222https://openstd.samr.gov.cn/bzgk/gb. This table consists of 6,763 Chinese characte rs, divided into 3,755 primary and 3,008 secondary characters. It contains most commonly used Chinese characters. In order to obtain the detailed information of this coding table, we collected various basic attributes of each Chinese word, such as its phonetic, radical, phonetic notation, strokes, outside strokes (the strokes except the radical), stroke order, structure, Unicode, Wubi code, Cangjie code, Zheng code, Four-corner code, as well as glyph images from different historical periods. Despite the large amount of data collected, we note issues with the quality and completeness of the data. To address these issues, we manually annotate the collected information to ensure a more accurate and comprehensive dataset. We give the attributes of the word 海 (sea) in the dataset as shown in Fig.3. In order to understand and better utilize these properties, we classify these features into the following three categories:
  • Phonetic Features:Phonetic which uses Latin letters to represent the pronunciation of Chinese characters. Phonetic notation is a phonetic system that uses symbols to represent the sounds of Chinese characters. Zheng Code is a Chinese character input method that assigns codes based on the initials of the phonetic pronunciation.
@@ 4.2 In Context Learning for CSC
A variety of studies Xie et al. (2022); Dai et al. (2023); Bansal et al. (2023) have revealed that LLMs exhibit exceptional in-context learning capabilities Dong et al. (2023). In-context learning can quickly improve the task-specific performance of LLMs by providing them with a limited set of task-related examples, which requires extremely high computational resources and is extremely expensive and difficult to fine-tune LLMs. In-context learning can quickly adapt to most LLMs without the need for separate training for each LLM. As mentioned in Section 3.3, when LLMs perform CSC tasks without specific task constraints on input prompts, the answers given by LLMs are very likely to be inconsistent with the existing evaluation indicators of CSC tasks. Therefore, in order to accurately and objectively explore the performance ability of LLMs on CSC tasks, we design the task-specific prompts as shown in Fig.2.
@@
First of all, we give the identity and task description of LLMs in the prompt, according to Li et al. (2023a), we know that giving LLMs emotional motivation (you are an excellent Chinese Spell Checking model) can effectively stimulate LLM’s comprehension. The task description requires the LLMs to only correct spelling mistakes, thus limiting the LLM’s semantic rephrasing of the input sentence. Since CSC task requires the input sentence and the output sentence to be equal in length, the LLMs is required not to add any other explanations and descriptions when performing CSC task, so as to require that the length of the sentence output by the LLMs and the input error correction sentence are equal. Please note that this does not mean that LLMs will not semantically rewrite the input sentence or change the sentence length after such a request. This can only minimize the role of LLMs in these issues.
@@
Next, We give three pairs of input sentences and their corrected sentences, carefully selected from the corresponding training set of the dataset we use. From Liu et al. (2010), Chinese text errors are primarily caused by characters that are visually and phonetically similar. These three sentence pairs contain a set of phonetic errors, a set of glyph errors, and a set of correct sentences that do not need to be corrected.
@@
We add specific task requirements to the input sentences of three pairs of sentences, which is different from the previous task description. We concretize the task here, asking LLMs to find spelling errors in sentences and replace wrong words with correct words. This allows LLMs to have a clearer understanding of CSC tasks. Next, we append the length of the input sentence to the end to indicate the LLMs that we limit the length of the output sentence in the first prompt. We also include the phonetic and radical information of each Chinese character in the input sentence, drawn from Chinese Rich Semantic Corpus outlined in Section 4.1. It is hoped that by adding this information to the prompt, the LLMs can strengthen their understanding of the input sentence. Therefore, LLMs can better correct phonetic and visual errors that may occur in the sentences. These are in the form of four pairs of historical dialogues that form a few-shot prompt input to LLMs, hoping to stimulate the in-context learning capabilities of LLMs at once.
For the sentences to be corrected in the test set, we use the prompt of the input sentence in the few-shot prompt to construct the same prompt for the sentences to be corrected in the testset and then enter LLMs to hope that the LLMs will correct the spelling of the error corrected sentences through in-context learning. Considering the length limit of LLMs on the length of the input (including historical dialogue), and the deterioration of the semantic understanding of LLMs with long input lengths. Before spell-checking each sentence, we clear the historical memory of the LLMs and add our few-shot prompt. In order to verify that the prompt we proposed is effective, we counted the number of sentences whose length remained unchanged before and after using a specific prompt and not using a specific prompt. The results are in Table 1. Experiments were performed using the API of GPT-3.5-Turbo.
Data without prompt
SIGHAN15 578 724
LAW 83 125
MED 237 262
ODW 294 347
Table 1: The number of sentences whose length does not change using specific prompts and not using specific prompts on different data sets
@@ 5.1 Datasets and Evaluation Metrics
Datasets: In the evaluation of CSC task, we use two highly used datasets. The first, SIGHAN15 Tseng et al. (2015), consists of handwritten samples from learners of Chinese as a Second Language (CSL). These samples provide a rich source of real-world language usage by non-native speakers, offering insights into common errors and patterns in learning Chinese. The second dataset Lv et al. (2023) is specialized and segmented into three distinct domains to provide a more comprehensive understanding of language use in specific contexts. For the LAW domain, data is sourced from the stems and choices of multiple-choice questions in judicial examinations, reflecting the formal and technical language of the legal field. The MED domain encompasses data from question-and-answer pairs drawn from online medical consultations, showcasing the specific terminology and communication style in healthcare. The ODW (Official Document Writing) domain includes data from various official documents such as news, policies, and state reports on national conditions, representing formal and structured writing styles. The statistics of the test data from these four datasets we used are shown in Table 2.
@@
Evaluation Metrics: Following the previous work Xu et al. (2021); Lv et al. (2023), we evaluate the performance of CSC at two levels: detection and correction. At the detection level, a sentence is considered correct if it successfully identifies all spelling errors. At the correction level, the model not only identifies but also rectifies all erroneous characters by replacing them with the correct ones. We provide accuracy, precision, recall, and F1 scores for both evaluation levels.
@@ 5.2 Baselines
In order to explore whether Chinese Rich Semantics in LLMs we proposed can be improved on CSC task, we choose two widely foundational models.
@@
glm-3-turbo [^4] https://open.bigmodel.cn/dev/api#glm-3-turbo: An open-source bilingual (Chinese-English) chat model. The glm-3-turbo is a variant of ChatGLM3 that is specifically optimized for dialogue generation tasks, and it has also achieved excellent results in various natural language processing tasks.
To evaluate the performance of LLMs, we choose advanced CSC models as our baseline references:
BERT Devlin et al. (2019): The basic BERT encodes the input sentence to get semantic information, followed by using a classifier to select the correct character from the vocabulary.
ChineseBERT Sun et al. (2021): The basic ChineseBERT encodes the input sentence to get semantic, phonetic, and graphical information, then use a classifier to select the correct character from the vocabulary.
ReaLiSe Xu et al. (2021): The CSC model captures semantic, phonetic, and graphical information of input characters using multimodality for prediction.
@@ 5.3 Implementation Details
In all CSC models in comparative experiments, the parameters are set to the optimal parameters in the open source code corresponding to the model. In order to achieve few-shot scenarios, all comparative experiments randomly selected 10 pieces of data from the training set, trained for 1000 rounds, and then selected the round with the best performance. The LLMs experiment uses the same few-shot prompt designed for CSC task as RS-LLMs, but does not add semantic information. In gpt-3.5-turbo-related experiments, we set the temperature to 0.8. In glm-3-turbo-related experiments, we set the temperature to the default 0.95. All experiments were conducted on a RTX-4090 with 24G memory.
@@
glm-3-turbo 45.6 28.9 40.1 33.6 42.2 23.9 33.1 27.7
RS-glm-3-turbo 46.4 29.7 40.1 34.2 43.5 25.3 34.2 29.1
gpt-3.5-turbo 37.0 24.1 38.6 29.7 35 21.6 34.6 26.6
RS-gpt-3.5-turbo 46.0 27.3 36.0 31.1 44.6 25.1 33.1 28.5
@@
glm-3-turbo 57.0 34.7 49.6 40.8 52.6 27.9 39.8 32.8
RS-glm-3-turbo 59.4 37.4 47.4 41.8 56.0 31.5 39.8 35.2
gpt-3.5-turbo 30.8 17.5 32.3 22.7 28.8 15.1 27.9 19.6
RS-gpt-3.5-turbo 39.0 20.7 34.1 25.8 36.4 17.2 28.3 21.4
LAW BERT 14.4 8.8 12.3 10.3 12.7 1.35 0.4 0.6
ChineseBERT 15.0 12.3 14.3 13.2 13.5 0.8 1.6 1.1
ReaLiSe 23.3 15.4 18.2 16.7 20.2 1.6 1.5 1.5
gpt-3.5-turbo 46.6 31.9 45.9 37.6 43.8 28.1 40.4 33.1
RS-gpt-3.5-turbo 56.2 37.6 43.9 40.5 54.0 33.9 39.6 36.5
@@
gpt-3.5-turbo 65.0 52.3 66.0 58.0 61.8 47.4 59.9 52.9
RS-gpt-3.5-turbo 75.6 63.7 62.2 62.9 73.4 58.4 58.0 58.7
Table 3: The performance of all baselines and RS-LLMs. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo and RS-glm-3-turbo means RS-LLM on glm-3-turbo. The glm-3-turbo and gpt-3.5-turbo utilizing identical few-shot prompts as RS-LLM, but without semantic information.
@@
Figure 5: Correct-F1 score trends on the test dataset.

### v3

@@ Abstract
Chinese Spell Checking (CSC) is a widely used technology, which plays a vital role in speech to text (STT) and optical character recognition (OCR). Most of the existing CSC approaches relying on BERT architecture achieve excellent performance. However, limited by the scale of the foundation model, BERT-based method does not work well in few-shot scenarios, showing certain limitations in practical applications. In this paper, we explore using an in-context learning method named RS-LLM (Rich Semantic based LLMs) to introduce large language models (LLMs) as the foundation model. Besides, we study the impact of introducing various Chinese rich semantic information in our framework. We found that by introducing a small number of specific Chinese rich semantic structures, LLMs achieve better performance than most of the BERT-based model on few-shot CSC task. Furthermore, we conduct experiments on multiple datasets, and the experimental results verified the superiority of our proposed framework.
{CJK}
UTF8gbsn
@@
Ming Dong1,2,3, Yujing Chen1,2,3, Miao Zhang4, Hao Sun1,2,3, Tingting He1,2,3,thanks:   Corresponding author.   1Hubei Provincial Key Laboratory of Artificial Intelligence and Smart Learning, 2National Language Resources Monitoring and Research Center for Network Media, 3School of Computer, Central China Normal University, Wuhan, China 4School of Computer Science and Information Engineering, Hubei University, Wuhan, China {dongming,haosun,tthe}@ccnu.edu.cn {yujingchen}@mails.ccnu.edu.cn {miaozhang}@hubu.edu.cn
@@ 1 Introduction
As an ideogram, the usage and structure of Chinese are very different from English, which leads to different challenges for CSC and SC. First of all, the pronunciation of Chinese varies greatly, and it is difficult to easily infer the glyphs from the pronunciation. When listening to a piece of Chinese speech, if you do not understand the context, the text result obtained based on the speech is likely to contain a large number of homophones. In addition, the glyph structure of Chinese is more diverse, resulting in more types of errors. Therefore, CSC task mainly needs to handle two types of error texts. One is a text based on the STT system, which contains a large number of homophonic errors. The other is text generated by OCR-based systems that mainly contains glyph errors. To address these two types of spelling errors, most of the existing studies use models based on BERT architecture, then introducing the external glyph-phonetic features Ji et al. (2021); Xu et al. (2021); Ji et al. (2021).
@@ 3.2 Task Definition
Formulate definition of CSC: CSC focuses on identifying and rectifying spelling errors in a given text sequence. Essentially, it involves processing a sequence of characters $X=\{x_{1},x_{2},...,x_{n}\}$ and producing a corrected sequence $Y=\{y_{1},y_{2},...,y_{n}\}$, where $n$ is the number of word. Here, $X$ denotes the initial text with potential errors, and $Y$ signifies the amended, correct text. The two sequences $X$ and $Y$ have the same length.
@@
3.3 Difficulties of LLMs based CSC Tasks
The versatility of LLMs gives them significant text polishing capabilities. Since there are no specific restrictions on input prompts, LLMs tend to perform freely in CSC tasks. However, free play of LLMs may result in LLMs outputting sentences that are completely grammatically correct. However, the sentences output by LLMs are different from the standards established by existing CSC datasets and evaluation indicators. Therefore, these existing traditional datasets are used without specific restrictions on input prompts. It becomes challenging to objectively and realistically evaluate the spell checking performance of LLMs. An example of using LLMs (ChatGPT) to perform a CSC task without setting specific restrictions on input prompts is shown in Fig.1. From Fig.1, it can be observed that two main problems are prone to occur in LLMs when performing CSC tasks, one is the length of the input sentence and the output sentence are inconsistent. For instance, the sentence "I am really bald (我真胡秃)", which should be corrected to "I am really confused (我真糊涂)". However, LLMs correct this sentence to "I am really sorry (我真的抱歉)". Another problem is that LLMs easily rewrite the input sentences semantically. For example, "I hope you have a good time dancing without (希望你们好好的跳无)", in which "dancing without (跳无)" should be corrected to "dancing (跳舞)". Instead, LLMs correct this sentence to "I hope you have a good New Year (希望你们好好的过年)".
Figure 2: Task-specific few-shot prompts for CSC tasks. We marked the semantic structure information (speech and radicals) and key information related to the task features in the prompts in different color.
@@ 4.1 Chinese Rich Semantic Corpus
Chinese, a logo-graphic language, inherently possesses a rich semantic depth in its character glyph, potentially enhancing the expressiveness of LLMs. Our work focuses on the GB2312 simplified Chinese coding table, a standard set by China’s State Administration of Standards on May 1, 1981222https://openstd.samr.gov.cn/bzgk/gb. This table consists of 6,763 Chinese characters, divided into 3,755 primary and 3,008 secondary characters. It contains the most commonly used Chinese characters. In order to obtain the detailed information of this coding table, we collected various basic attributes of each Chinese word, such as its phonetic, radical, phonetic notation, strokes, outside strokes (the strokes except the radical), stroke order, structure, Unicode, Wubi code, Cangjie code, Zheng code, Four-corner code, as well as glyph images from different historical periods. Despite the large amount of data collected, we note issues with the quality and completeness of the data. To address these issues, we manually annotate the collected information to ensure a more accurate and comprehensive dataset. We give the attributes of the word 海 (sea) in the dataset as shown in Fig.3. In order to understand and better utilize these properties, we classify these features into the following three categories:
  • Phonetic Features:Phonetic uses Latin letters to represent the pronunciation of Chinese characters. Phonetic notation is a phonetic system that uses symbols to represent the speech of Chinese characters. Zheng Code is a Chinese character input method that assigns codes based on the initials of the phonetic pronunciation.
@@
We publish the information we collect at dropbox.  [^3] https://www.dropbox.com/scl/fo/0r1jw4l1ex3w0lyojsfpf/AGVMdlOFpwDqoOFrxqWXJko?rlkey=18wke1vmhj6muwufvu3aazji2&st=vgh1drbi&dl=0
@@ 4.2.1 Motivation
A variety of studies Xie et al. (2022); Dai et al. (2023); Bansal et al. (2023) have revealed that LLMs exhibit exceptional in-context learning capabilities Dong et al. (2023). In-context learning can quickly improve the task-specific performance of LLMs by providing a limited set of task-related examples, which can quickly adapt to most LLMs without the need for separate training for each LLM. As mentioned in Section 3.3, when LLMs perform CSC tasks without specific task constraints on input prompts, the answers generated by LLMs are very likely to be inconsistent with the existing evaluation indicators of CSC tasks. Therefore, in order to accurately and objectively explore the performance ability of LLMs on CSC tasks, we design the task-specific prompts as shown in Fig.2.
4.2.2 Prompt Design
@@
First of all, we give the identity and task description in the prompt. According to Li et al. (2023a), we know that Role Attribution (“You are an excellent Chinese Spell Checking model") can effectively stimulate LLMs’ comprehension. The task description requires the LLMs to only correct spelling mistakes, thus limiting the LLMs’ semantic rephrasing of the input sentence. Since the evaluation of CSC task requires the input sentence and the output sentence to be same length, the LLMs are required not to add any other explanations and descriptions of output, so as to ensure the length remains unchanged.
@@
Then, we give three pairs of input sentences and their corrected sentences as examples, carefully selected from the corresponding training set of the dataset we use. From Liu et al. (2010), Chinese text errors are primarily caused by characters that are visually and phonetically similar. These three sentence pairs contain a set of phonetic errors, a set of glyph errors, and a set of correct sentences that do not need to be corrected.
@@
Next, we add specific task requirements to the input sentences of three pairs of sentences, which is different from the foremost task description. The foremost task description requires LLMs to correct the spelling errors of the input sentences. We concretize the task here, asking LLMs to find spelling errors in sentences and replace wrong word with correct one. This allows LLMs to have a clearer comprehension of CSC task. We append the length of the input sentence to the end to indicate the LLMs that we limit the length of the output sentence in the first prompt. We also include the phonetic and radical information of each Chinese character in the input sentence, drawn from Chinese Rich Semantic Corpus outlined in Section 4.1. It is hoped that by adding this information to the prompt, the LLMs can strengthen their understanding of the input sentence. Therefore, LLMs can better correct phonetic and visual errors that may occur in the sentences. These are in the form of four pairs of historical dialogues that form a few-shot prompt input to LLMs, hoping to stimulate the in-context learning capabilities of LLMs at once.
Considering the length limit of LLMs on the length of the input (including historical dialogue), and the deterioration of the semantic understanding of LLMs with long input lengths. For each sentence to be corrected in the test set, we clear the historical memory of the LLMs and add our few-shot prompt.
4.3 Introspection Mechanism
A significant challenge with using LLMs for spell checking is that LLMs tend to over-modify and arbitrarily change sentence lengths. In order to avoid the impact of this change on the evaluation indicators, we use a Introspection mechanism.
Specifically, after LLMs generate the correction sentence, we send this answer and the original input sentence to LLMs again. LLMs are required to introspect two questions: 1) Whether the lengths of the two sentences are consistent. 2) Whether the added rich semantic information is effectively used in this error correction process. Only when the answers to both questions are "yes" will the answer be output as the final error correction result. Otherwise, the current conversation will be added to the historical conversation and provided as context to the model. In the conversation, it indicates that the sentence length in this answer does not match the input sentence length or the semantic information is not used, and then sends a reply request to LLMs again. In the introspection mechanism of this experiment, we set the maximum number of loops to 5. If the answer cannot be obtained after five requests to LLMs, it is judged that LLMs cannot correct the sentence and uses original input sentence without introspection as the answer. And consider that the model’s ability to understand long context will deteriorate. We only add the latest round of dialogue in each loop within the context of the original design.
@@ 5.1 Datasets and Evaluation Metrics
Datasets: To evaluate the effectiveness of our proposed method, we choose two widely used datasets. The first, SIGHAN15 Tseng et al. (2015), consists of handwritten samples from learners of Chinese as a Second Language (CSL). These samples provide a rich source of real-world language usage by non-native speakers, offering insights into common errors and patterns in learning Chinese. The second dataset Lv et al. (2023) is specialized and segmented into three distinct domains to provide a more comprehensive understanding of language use in specific contexts. For the LAW domain, data is sourced from the stems and choices of multiple-choice questions in judicial examinations, reflecting the formal and technical language of the legal field. The MED domain encompasses data from question-and-answer pairs drawn from online medical consultations, showcasing the specific terminology and communication style in healthcare. The ODW (Official Document Writing) domain includes data from various official documents such as news, policies, and state reports on national conditions, representing formal and structured writing styles. The statistics of the test data from these four datasets we used are shown in Table 1.
@@
Evaluation Metrics: By following the existing work Xu et al. (2021); Lv et al. (2023), we evaluate the performance on two metrics: detection and correction. For detection, a sentence is considered correct if it successfully identifies all spelling errors. For correction, the model not only identifies but also rectifies all erroneous characters by replacing them with the correct ones. We provide accuracy, precision, recall, and F1-scores for both metrics.
@@ 5.2 Baselines
We choose two widely used LLMs as foundational models.
@@
ChatGLM2-6B [^5] https://huggingface.co/THUDM/chatglm2-6b: An open-source bilingual (Chinese-English) chat model. ChatGLM2-6B employs a GLM-based Du et al. (2022) hybrid objective function and has been pre-trained on 1.4 trillion bilingual tokens and human preference alignment training.
We choose several advanced CSC models as baselines:
BERT Devlin et al. (2019): BERT encodes the input sentence to get semantic information, followed by using a classifier to select the correct character from the vocabulary.
ChineseBERT Sun et al. (2021): ChineseBERT encodes the input sentence to get semantic, phonetic, and graphical information, then use a classifier to select the correct character from the vocabulary.
ReaLiSe Xu et al. (2021): This CSC model captures semantic, phonetic, and graphical information of input characters using multimodality for prediction.
Scope Li et al. (2022): This CSC model introduces an auxiliary task of Chinese pronunciation prediction (CPP) to improve CSC task.
@@ 5.3 Implementation Details
For all methods, the settings of hyper-parameters follow the optimal parameters in the open source code corresponding to the model. In order to achieve few-shot scenarios, all contrast experiments randomly selected 10 samples of data from the training set, trained for 1000 epochs, and then selected the best performance. The LLMs experiment uses the same few-shot prompt designed for CSC task as RS-LLMs, but does not add semantic information. All experiments were conducted on a RTX-4090 with 24G memory.
Data without prompt with prompt
SIGHAN15 578 724
LAW 83 125
MED 237 262
ODW 294 347
Table 2: The number of sentences whose length does not change using specific prompts and not using specific prompts on different data sets. Experimental result statistics are based on gpt-3.5-turbo.
@@
Scope 64.7 61.7 34.0 43.9 58.1 37.2 20.5 26.5
ChatGLM2-6B 18.3 7.1 10.1 8.3$\pm$1.1 17.0 3.3 5.3 4.1$\pm$2.8
RS-ChatGLM2-6B 30.7 14.9 18.0 16.3$\pm$2.3 28.6 12.4 15.6 13.9$\pm$3.9
gpt-3.5-turbo 36.4 22.3 33.0 26.6$\pm$3.1 34.3 19.3 28.6 23.1$\pm$3.5
RS-gpt-3.5-turbo 50.6 32.5 41.6 36.5$\pm$4.7 48.1 31.2 40.8 35.4$\pm$5.9
LAW BERT 14.4 8.8 12.3 10.3 12.7 1.35 0.4 0.6
ChineseBERT 15.0 12.3 14.3 13.2 13.5 0.8 1.6 1.1
ReaLiSe 23.3 15.4 18.2 16.7 20.2 1.6 1.5 1.5
Scope 66.2 50.4 48.6 49.5 58.6 35.0 33.7 34.3
ChatGLM2-6B 36.6 18.5 25.5 21.4$\pm$2.7 34.8 15.9 21.9 18.5$\pm$3.3
RS-ChatGLM2-6B 45.2 24.2 25.1 24.6$\pm$3.4 40.4 22.7 24.8 23.7$\pm$3.2
gpt-3.5-turbo 48.8 30.2 38.8 34.0$\pm$3.6 46.2 26.2 33.7 29.5$\pm$3.6
RS-gpt-3.5-turbo 64.6 46.6 56.1 50.9$\pm$9.4 60.8 40.4 48.6 44.1$\pm$8.7
@@
Scope 66.4 45.6 53.9 49.4 56.5 27.3 32.3 29.6
ChatGLM2-6B 31.2 15.7 23.0 18.7$\pm$5.6 28.4 14.7 19.2 16.7$\pm$6.5
RS-ChatGLM2-6B 45.2 30.4 33.3 31.8$\pm$6.7 39.9 22.6 30.6 26.0$\pm$4.2
gpt-3.5-turbo 41.4 16.8 30.1 21.5$\pm$1.2 38.6 13.3 23.9 17.1$\pm$2.7
RS-gpt-3.5-turbo 56.0 31.0 45.2 36.8$\pm$8.7 43.3 25.3 39.3 30.8$\pm$6.8
@@
Scope 75.0 65.5 58.7 62.0 70.8 56.5 50.7 53.5
ChatGLM2-6B 47.6 28.6 32.5 30.4$\pm$6.5 44.2 27.0 31.1 28.9$\pm$7.8
RS-ChatGLM2-6B 56.4 32.1 37.8 34.7$\pm$2.7 50.8 29.8 32.4 31.0$\pm$3.2
gpt-3.5-turbo 63.0 45.8 50.0 47.8$\pm$8.5 59.2 39.2 42.8 40.9$\pm$7.9
RS-gpt-3.5-turbo 72.4 59.1 64.9 61.9$\pm$2.6 70.2 50.2 57.6 53.6$\pm$5.1
Table 3: The performance of all baselines and RS-LLMs. RS-gpt-3.5-turbo means RS-LLM on gpt-3.5-turbo and RS-ChatGLM2-6B means RS-LLM on ChatGLM2-6B. ChatGLM2-6B and gpt-3.5-turbo only utilize identical few-shot prompts as RS-LLM, without semantic prompt and introspection.The bold information indicates the best results except Scope (Scope is the best BERT-based model in few-shot scene), and $\pm$ indicates the error band of the results.
@@
Figure 5: The impact of different number of prompt samples.